Plot P=f(M) plotted as-is, a plot that delivers a straight line, and the quantitative relationship (function) between the two variables.

P

M

4.00

0.08

4.40

0.25

4.80

0.40

5.32

1.00

5.67

2.00

5.90

3.00

6.09

4.00

6.27

5.00

6.45

6.00

6.64

7.00

6.87

8.00

7.22

9.00

7.55

9.50

8.27

9.80

8.70

10.00
Answer #1

Sum of X = 75.03
Sum of Y = 94.15
Mean X = 5.002
Mean Y = 6.2767
Sum of squares (SSX) = 196.2188
Sum of products (SP) = 67.3627

Regression Equation = ŷ = bX + a

b = SP/SSX = 67.36/196.22 = 0.3433

a = MY - bMX = 6.28 - (0.34*5) = 4.55946

ŷ = 0.34X + 4.56